INTRODUCTION Jaden McCoy operated a dairy goat farm in Soddy-Daisy, Tennessee, and was considering expanding the rental of hiMcCoy believed his success with the initial goat rental job would lead to additional land clearing work at the resort and - vthe shepherd pointed out the truck McCoy used for the resort job did not have the power, suspension, or drive train needed toAs McCoy began to put his numbers together, he realized the larger trailer would enable him to pursue more jobs each year thaExhibit 1 ACTUAL REVENUES AND COSTS FOR THE RESORT LAND-CLEARING JOB in U.S. dollars) Per day Project $375.00 $2,625.00 CalenExhibit 2 FIXED AND VARIABLE COSTS FOR RESORT JOB (in U.S. dollars) Description Truck and trailer operation and maintenance cQuestion 1: Compute the number of calendar days needed to complete a one-acre job using the current truck/trailer combination
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Note: I have answered part 1 of the question 1 as the solution is too lenghty and HOMEWORKLIB RULES restricts it to a constrained leve. Be sure to ask the rest of the questions in a seperate part, I will be more than happy and willing to help !!!

There are certain assumptions we must make before proceeding to the case:

  • The pace of deterioration and the hour of procurement isn't given in the inquiry. In this way, we accept that the expense of bigger trucks and trailers has just been borne by the organization and no installments will be made towards them for example no costs will be brought about in getting them.
  • The expense for the present trucks and trailers is additionally not given, so we expect that no installments as well as expenses are to be incorporated while making counts.

According to question, we have to ascertain the quantity of schedule days expected to finish the activity utilizing current trucks/trailers and the bigger trucks/trailers. At the end of the day, we have to ascertain the Break Even Days when the in any event the absolute expense caused in the activity will be recuperated.

In this way, we liken the income and the complete expenses in the two situations..

Case 1: Current Trucks/ Trailers

Let the number of calendar days required be x.

So, Total Revenue = 375x

Total Fixed Cost = Fencing Cost + Set up Cost = 0.75*962 + 100

Total Variable Cost/ Day = Shepherd and Dogs Cost + Truck Maintenance & Operations Costs

= >190 + 0.63*80 (Data taken from Exhibit 2) = 240.40

For breakeven, Revenue = Costs

=> 375x = 240.4x + 821.5....................(1)

On solving the above equation, we get, x = 6.103 days

which is not possible, so we round it off to 7 days

But, to calculate the max number of jobs that can be taken with this combination of Trucks/ Trailers, we use x = 6.103.

Number of Jobs/ year = 365/ 6.103 = 59.8. We round it off to 59 as it cannot be 60

Case 2: Larger Trucks/ Trailers

Considering Exhibit 4,

Equating revenue = costs,

375y = (1.25*80 + 190)y + 100 + 0.75*962 ( y is the days)

On solving the above equation, we get, y = 9.66 days, Rounding it off to 10

Number of Jobs/ year = 365/ 9.66 37.78 or 37 jobs/ year.